OpenStudy (anonymous):

Celeste wants to have her hair cut and permed and also go to lunch. She knows she will need $86. The perm cost twice as much as her haircut and she needs $5 for lunch. How much does the perm cost ?

OpenStudy (fibonaccichick666):

Do you know how to set up a system of equations?

OpenStudy (anonymous):

I do not can you explain ?

OpenStudy (fibonaccichick666):

Ok, so what could we add up to get that total $86?

OpenStudy (anonymous):

ok so the $5 for lunch

OpenStudy (fibonaccichick666):

ok, what else? Let's say the perm costs x dollars

OpenStudy (anonymous):

the perm cost twice as much as her haircut

OpenStudy (fibonaccichick666):

right, so if the *haircut* costs x dollars, the perm costs? (sorry had a dyslexic moment)

OpenStudy (anonymous):

i have no idea i've never set up this system you'd multiply 5 by ?

OpenStudy (fibonaccichick666):

You said "the perm costs twice as much as her haircut" ^ ^ ? x

OpenStudy (anonymous):

5*2?

OpenStudy (fibonaccichick666):

you don't do anything to the 5

OpenStudy (anonymous):

2*5

OpenStudy (fibonaccichick666):

the 5, is a five, is a five. It does not affect the perm

OpenStudy (fibonaccichick666):

so you add 3 things together to get $86 right?

OpenStudy (anonymous):

so so we divide 86 ?

OpenStudy (fibonaccichick666):

What 3 things in terms of x?

OpenStudy (anonymous):

so thre's 81 dollars left so is estimated around 40 dollars divided evenly each ?

OpenStudy (fibonaccichick666):

no, you don't divide we are just setting up the eq

OpenStudy (fibonaccichick666):

you want something that looks like this a+b+c=86

OpenStudy (fibonaccichick666):

no leave the 5 alone

OpenStudy (anonymous):

ohh ok so do we plug number in any numbers or ?

OpenStudy (fibonaccichick666):

yes, ok smaller ex. you go to get gas and a coffee. The gas is 4 times the price of the coffee. The total was $25

OpenStudy (fibonaccichick666):

So call coffee c, this yeilds the eq: 4c+c=25 Do you see how I got that?

OpenStudy (anonymous):

ohh ok so your just using the coffee to find the solution ?

OpenStudy (fibonaccichick666):

yes, now let's say I got a coffee a candy bar and gas. The candy bar was 1.25 now i'm spending 26.25

OpenStudy (anonymous):

so that adds a variable to your equation ?

OpenStudy (fibonaccichick666):

that gives us, 4c+c+1.25=26.25

OpenStudy (fibonaccichick666):

nope, we know the price, so, it's just there

OpenStudy (fibonaccichick666):

Do you see now?

OpenStudy (anonymous):

oh ok and yea so how would you figure that out when all we know is that it cost 25 dollars and the gas is four time the price of the coffee

OpenStudy (fibonaccichick666):

Well see the eq?

OpenStudy (fibonaccichick666):

we have 4c+c+1.25=26.25. (I changed the final price to make things easier. So do we have any like terms we can combine?

OpenStudy (anonymous):

yea we have 2 c's so do you combine them like so \[c ^{2}\]

OpenStudy (fibonaccichick666):

no, they are not being multiplied. Just add them

OpenStudy (anonymous):

so give me an example of what it'd look like ?

OpenStudy (fibonaccichick666):

what is 3+4?

OpenStudy (anonymous):

the perm cost 30 bucks

OpenStudy (fibonaccichick666):

idk, I haven't solved it yet

OpenStudy (fibonaccichick666):

but, no that is not correct

OpenStudy (anonymous):

so it'd be c+c

OpenStudy (fibonaccichick666):

first, what is 3+4?

OpenStudy (fibonaccichick666):

what is (3+4)c? What is 3c+4c?